A cylindrical can of orange concentrate has a diameter of 6 cm and a height of 12 cm. The instruction on the can are to mix the orange with 3 cans of water.
a) Will the mixture fit into a cylindrical jug with diameter of 10cm and height 20cm?
b) How high would a jug with diameter 10cm have to be, to hold the mixture?

Answer by Cromlix(4381)

Volume of concentrate can = Pi*r^2*h
= Pi*3^2*12
= 339.29cm^3
Volume of three concentrate cans = 1017.88cm^3
Total = 1357.17cm^3
Volume of cylindrical jug = Pi*r^2*h
= Pi*5^2*20
= 1570.8cm^3
a)Therefore, yes the concentrate + water will fit
in the jug.
b)The total quantity of liquid = 1357.17cm^3
Pi* r^2* h = 1357.17cm^3
Pi* 5^2* h = 1357.17cm63
h = 1357.17cm^3/Pi * 25
h = 17.3cm.
Therefore this is the height required to
hold the mixture in a jug with diameter 10cm.
Hope this helps.
:-)